DMP(1) DMP(1) NAME dmp−translateTroffoutputtolow-levelMetaPostinput SYNOPSIS dmp[option]ditrofffile[mpxfile] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. ThedmpprogramgeneratesMetaPostsourcecodefromtheditrofffile.Thegeneratedcodeis writteninmpxfileifitisspecified,andtostandardoutputotherwise. dmpistypicallycalledfromscriptslikemakempx,notfromthecommandline. OPTIONS dmpacceptsthefollowingoptions: --helpPrinthelpmessageandexit. --version Printversioninformationandexit. SEEALSO mpost(1),makempx(1). Web2C7.32x 4January1998 1